OpenStudy (anonymous):

what is 5/8 of 52 miles?

OpenStudy (radar):

\[5/8 \times52/1=5/2\times13/1=65/2=32.5mi\]

OpenStudy (anonymous):

thank you.

OpenStudy (amistre64):

too far, cant walk it... aint got the gas to drive it.... and when you get there everyones gone :)